OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 2 sin^2(x) = sin (x)

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

what values of x would make sin(x) = 0? Since (2) (0)^2 = 0 0= 0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so would it be 0,pi,pi/6,5pi/6?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

I'd say 0, pi, and 2pi

OpenStudy (anonymous):

sin pi/6 = 0.5 and sin 5pi/6 = 0.5 as well. Refer to circle of cos and sin here: http://fac-web.spsu.edu/math/edwards/1113/unit2.gif

OpenStudy (anonymous):

ok can you help me with one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Find all solutions to the equation. sin2x + sin x = 0 step by step would be fantastic

OpenStudy (anonymous):

Just to make sure it's sin (2x) right? not \[\sin ^{2} x\]

OpenStudy (anonymous):

oh sorry its the second one sin^2(x)

OpenStudy (anonymous):

again, sin^2 x = (sin(x))^2 0+ 0 = 0, so x = 0, pi, 2pi are answers

OpenStudy (anonymous):

there are more though

OpenStudy (anonymous):

what are the others

OpenStudy (anonymous):

(-1)^2 + (-1) = 0 so any sin(x) = -1 are answers. Find those values

OpenStudy (anonymous):

how do i find those values

OpenStudy (anonymous):

the circle I linked you. (cos, sin) sin is the second value

OpenStudy (anonymous):

it says 3pi/2

OpenStudy (anonymous):

Yes, so x= 0, pi, 3pi/2, and 2pi

OpenStudy (anonymous):

ok is there a referance video or something i can see to show me how to do that?

OpenStudy (anonymous):

to do what? find sin(x) = certain answers or solving this problem?

OpenStudy (anonymous):

solving this problem and others like it

OpenStudy (anonymous):

Then my best suggestion is your textbook. The videos are really broad on trigonometry so I doubt you can find one that teaches you to solve certain problems like these. They cover more concepts.

OpenStudy (anonymous):

If you have any questions, I'm here to help though. What part are you confused about?

OpenStudy (anonymous):

alright thank you so much

OpenStudy (anonymous):

Welcome :)